# A family of log-correlated Gaussian processes

### v1

@@ 1.2. A central limit theorem
One may ask whether the stochastic-integral representations and the convergence above hold for all functions $f$ such that $\iint f(x)f(y)\Gamma(x,y){\rm d}x{\rm d}y<\infty$. We impose the condition in ${\mathcal{F}}_{0}$ as it is convenient for establishing our limit theorems. Also, ${\mathcal{F}}_{0}$ include all smooth functions that decay rapidly at infinity. Lévy’s continuity theorem yields immediately the convergence of $G_{n}$ in the space of tempered distributions. For the generalized process $G$ indexed by ${\mathsf{M}}$, Lévy’s continuity theorem applies as long as the space of functions on ${\mathsf{M}}$ is a nuclear space [12, 33]. We do not pursue to strengthen the convergence in the space of distributions here.
@@
[…]
1.3. Further comments
We conclude the introduction with remarks on a few related results in the literature.
Remark 1.7.
In an inspiring paper, Neuman and Rosenbaum, [36] considered to characterize the limit process of fractional Brownian motions as $H\downarrow 0$. Therein, the authors investigated
provided an argument on why a shift also makes sense in practice in finance applications, and then investigated the limit. This limit theorem has been extended by Hager and Neuman, [18] to ${\mathbb{R}^{d}}$-indexed random fields, with a generalized fractional Brownian field replacing ${\mathbb{B}}^{H}$. One of the highlights of [18] is that, in order to obtain a non-degenerate log-correlated Gaussian processes, instead of working with a specific shift as in (1.10) or in our setup, a general set of assumptions are provided. Note that once a generalized Gaussian process is obtained in the scaling limit, a corresponding Gaussian multiplicative chaos can be introduced, and this has been discussed in [36, 18]. In fact, the focus of [18] is to establish the convergence of an exponential functional of process $X_{H}$ in (1.10) to the Gaussian multiplicative chaos directly without proving the convergence to the generalized processes.
@@ 1.3. Further comments
$\log\left(\frac{\sqrt{t}+\sqrt{s}}{\sqrt{t}-\sqrt{s}}\right)\mbox{ for all }0\leq s<t<\infty.$
Their limit theorem is in the same spirit as [36].
Remark 1.9.
The limit process $G$ introduced here can already be read from recent developments by Iksanov and Kotelnikova, [22] on the Karlin model, and it is from their results we first had the guess of the expression of $G$. The Karlin model is an infinite urn model of which the law of the sampling over $\mathbb{N}$ decays as $j^{-1/{2H}},H\in(0,1/2)$ as $j\to\infty$. The model was introduced by Karlin, [28], who showed that the partial sum of so-called odd-occupancy process (the counting process of urns occupied by an odd number of balls), normalized by $n^{H}$, scaled to a Gaussian random variable. The consideration of odd-occupancy process was inspired by Spitzer, [47] to modeling the on/off events for lightbulbs. The functional central limit theorem for the odd-occupancy process was established in [10] (strictly speaking on a variation of the original model), and the limit is a fractional Brownian motion with Hurst index $H$. This line of research has been extended to stable processes indexed general metric spaces [9, 14].
The Karlin model with $H=0$ has been recently investigated by Iksanov et al., [21] and Iksanov and Kotelnikova, [22]. This is the case where the sampling frequencies $(p_{j})_{j\in\mathbb{N}}$ satisfy, for $\rho(t):=|\{k\in\mathbb{N}:p_{k}\geq 1/t\}|$, that there exists a slowly varying function $g$ at infinity with $g(x)\to\infty$ as $x\to\infty$ such that $\lim_{x\to\infty}(\rho(\lambda t)-\rho(t))/g(t)=\log\lambda$ for all $\lambda>0$. Limit theorems have been established in this regime with new Gaussian processes in the limit in the aforementioned papers. Their representation is not immediately recognizable to be related to ours and we provide some details in Appendix A. It should be expected that from their model a counterpart of our Theorem 1.4 can be established (they examined $j$-occupancy processes instead of odd-occupancy processes). Since the analysis of Karlin model in general is more involved than Enriquez model, and yet they have the same scaling limit, we choose to present a central limit theorem on a variation of the latter model.
Remark 1.10.
Another example of fractional Brownian motions with Hurst index $H=0$ has been seen in the studies of random matrix theory [15, 29]. Their example is a centered self-similar Gaussian process with stationary increments.
The paper is organized as follows. Section 2 presents a stochastic-integral representation of the log-correlated Gaussian process $G(f)$. Section 3 provides a proof of Theorem 1.4. Section 4 provides the general framework for log-correlated Gaussian processes indexed by metric spaces, and also a central limit theorem based on a different aggregated model.
@@ Acknowledgements
Y.W. thanks Alexander Iksanov, Joseph Najnudel, and Yi Shen for several discussions. Y.W. was partially supported by Simons Foundation (MP-TSM-00002359), and a Taft Center Fellowship (2024–2025) from Taft Research Center at University of Cincinnati.
@@ 2. A stochastic-integral representation
We first recall the stochastic-integral representations of fractional Brownian motions and its decomposition (1.5) recently discovered in the studies of Karlin model [10, 9, 14]. Our reference for stochastic integrals with respect to stable random measures is Samorodnitsky and Taqqu, [44]: if $(E,{\mathcal{E}},\mu)$ is a measure space, $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on $(E,{\mathcal{E}})$ with control measure $\mu$, then the stochastic integral $\int f{\rm d}M_{\alpha}$ with control measure $\mu$ is defined as a symmetric $\alpha$-stable random variable with
@@
Let $(\Omega,{\mathcal{F}},\mathbb{P})$ be the by-default probability space that the random measures are defined on. Set $E=\mathbb{R}_{+}\times\Omega^{\prime}$ where $(\Omega^{\prime},{\mathcal{F}}^{\prime},\mathbb{P}^{\prime})$ is another probability space on which a standard Poisson process $N^{\prime}$ is defined, and $M_{2,H}$ is a Gaussian random measure (defined on $(\Omega,{\mathcal{F}},\mathbb{P})$) on $E=\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $C_{2H}r^{-2H-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ with $C_{\beta}=\beta 2^{-\beta}/\Gamma(1-\beta),\beta\in(0,1)$. Then, the following three Gaussian processes satisfy the decomposition (1.5):
${\mathbb{B}}_{t}^{H}$ :=+×$:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}{{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}M_{2,H}({\rm d}r{\rm d}\omega^{\prime}),$
$G^{H}_{t}$ :=+×$:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)M_{2,H}({\rm d}r{\rm d}\omega^{\prime}),$
$W^{H}_{t}$ :=+×$:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\widetilde{p}_{r}(t)M_{2,H}({\rm d}r{\rm d}\omega^{\prime}),$
with
@@
In this way, $G^{H}_{t}={\mathbb{B}}^{H}_{t}-W^{H}_{t}$ almost surely and one can view it as a fractional Brownian motion with a random shift. Note that ${\mathbb{B}}^{H}$ and $W^{H}$ are dependent by construction. These stochastic-integral representations were discovered in [9]; see also [14, Proposition 3.7]. Processes with similar representations as $G_{t}^{H}$ or ${\mathbb{B}}_{t}^{H}$ (the integrand being an indicator function possibly centered) often arise as scaling limits of counting processes associated to models with certain renewal structures; see (A.2) for the limit process of the $j$-occupancy process of the Karlin model with $H=0$, and [1] for a limit theorem on the counting process for a family of decoupled renewal processes.
We next introduce the following:
@@
where $M_{2}$ is a Gaussian random measure on +×$\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ and $f$ is from the following class:
(2.3) ${\mathcal{F}}_{0}:=\left\{f\in L^{\infty}([0,\infty)):\int_{0}^{\infty}s^{\delta}|f(s)|{\rm d}s<\infty\mbox{ for some $\delta>0$}\right\}.$
In particular, $f\in{\mathcal{F}}_{0}$ implies that ${\left\|f\right\|}_{1}<\infty$.
@@
For all $f\in{\mathcal{F}}_{0}$,
$\int_{0}^{\infty}\int_{0}^{\infty}|f(s)f(t)|\log\frac{s+t}{|t-s|}{\rm d}s{\rm d}t<\infty.$
As a consequence, the collection $\{G(f)\}_{f\in{\mathcal{F}}_{0}}$ is a family of centered Gaussian random variables that satisfy
(2.4) $G(a_{1}f_{1}+a_{2}f_{2})=a_{1}G(f_{1})+a_{2}G(f_{2})\mbox{ almost surely, for all }a_{1},a_{2}\in\mathbb{R},f_{1},f_{2}\in{\mathcal{F}}_{0},$
and also with covariance function (1.7).
@@
We use the identity (due to Frullani integral)
(2.5) $\int_{0}^{\infty}\frac{e^{-r(t-s)}-e^{-r(t+s)}}{r}{\rm d}r=\log\left(\frac{t+s}{t-s}\right),\quad 0\leq s<t.$
Without loss of generality, assuming $f\geq 0$. Then, it is equivalent to show
(2.6) $\int_{0}^{\infty}\int_{0}^{\infty}\int_{s}^{\infty}f(s)f(t)\left(e^{-r(t-s)}-e^{-r(t+s)}\right){\rm d}t{\rm d}s\frac{{\rm d}r}{r}\equiv\int_{0}^{\infty}\Psi_{f}(r)\frac{{\rm d}r}{r}<\infty.$
(We applied Fubini’s theorem since the integrand is non-negative.) We shall show
(2.7) $\Psi_{f}(r)=\int_{0}^{\infty}\int_{s}^{\infty}f(s)f(t)\left(e^{-r(t-s)}-e^{-r(t+s)}\right){\rm d}t{\rm d}s\leq C\left(r^{\delta}\wedge r^{-1}\right),\mbox{ for all }r>0,$
for some constant $C$ depending on $f\in{\mathcal{F}}_{0}$, whence (2.6) holds. By a change of variables write first
First, for all $r>0$ we have
and
We have proved (2.7) and the first part of the lemma.
For the second part, we first claim that
@@
We have just seen that the last expression is finite. Provided that the equality in (2.8) holds, then the first equality above follows from properties of stochastic integrals. So it suffices to prove (2.8), which follows from a careful application of Fubini’s theorem. Indeed, for all $0\leq s<t$,
@@
Here in the first step we applied Fubini’s theorem to interchange the two inner integrals (in this step we verify the condition in Fubini’s theorem for each $r>0$, as it simply holds that $\iint f(s)f(t)\displaystyle{\rm d}t={\left\|f\right\|}_{1}^{2}<\infty$), and in the second step we applied (2.9). In the last expression above, we can interchange the two integrals again by Fubini’s theorem, which, in combination of Frullani integral (2.5), leads to (2.8).
@@
where we follow the same notations for stochastic integrals earlier. Note that by stochastic Fubini theorem [37, Proposition 5.13.1], we have the representation
(2.11) $G^{(\epsilon)}(f)=\int_{0}^{\infty}\widetilde{G}^{(\epsilon)}_{t}f(t){\rm d}t,\mbox{ almost surely,}$,
@@ 2.1. An extension to a generalized stable process
$G_{\alpha}(f):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}tM_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),f\in{\mathcal{F}}_{0},$
where $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on $[0,\infty)\times\Omega^{\prime}$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$. So $G(f)=G_{2}(f)$ and we again have $G_{\alpha}(a_{1}f_{1}+a_{2}f_{2})=a_{1}G_{\alpha}(f_{1})+a_{2}G_{\alpha}(f_{2})$ almost surely for $a_{1},a_{2}\in\mathbb{R},f_{1},f_{2}\in{\mathcal{F}}_{0}$, in view of the following characterization. Recall that $M_{\alpha}$ is a symmetric $\alpha$-stable random measure and the corresponding stochastic integrals are random variables characterized by (2.1).
@@
For all $f\in{\mathcal{F}}_{0}$, $G_{\alpha}(f)$ is a symmetric $\alpha$-stable random variable with characteristic function:
${\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right)=\exp\left(-2\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}\right).$
@@
(2.13) +×$\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left|\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}{\rm d}\mathbb{P}^{\prime}\\ =\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}<\infty.$
By Hölder’s inequality and then an earlier estimate (2.7), we have

### v3

@@ 1.2. Related results in the literature
In an inspiring paper, Neuman and Rosenbaum, [37] considered to characterize the limit process of fractional Brownian motions, denoted by ${\mathbb{B}}^{H}$, as $H\downarrow 0$. Therein, the authors investigated
provided an argument on why a shift also makes sense in practice in finance applications, and then showed the limit of $X_{H}$ is a log-correlated Gaussian process. Later on, Hager and Neuman, [18] showed that the random shift $t^{-1}\int_{0}^{t}{\mathbb{B}}_{s}^{H}{\rm d}s$ can be relaced by a larger family of processes and the limits remain to be log-correlated (and also they considered $\mathbb{R}^{d}$-indexed random fields for $d\in\mathbb{N}$). Our process corresponds to a specific choice of the shift, essentially related to the decomposition of a fractional Brownian motion by a bi-fractional Brownian motion and a smooth Gaussian process due to Lei and Nualart, [30]. We provide more details in Remark 2.6.
@@
The limit process $G^{1/2}$ introduced here can already be read from recent developments by Iksanov and Kotelnikova, [22] on the Karlin model [28, 17], and it is from their results we first had the guess of the expression of $G^{1/2}$. The Karlin model is an infinite urn model of which the law of the sampling over $\mathbb{N}$ decays as $j^{-1/{2H}},H\in(0,1/2)$ as $j\to\infty$. The model was introduced by Karlin, [28], who showed that the partial sum of so-called odd-occupancy process (the counting process of urns occupied by an odd number of balls), normalized by $n^{H}$, scales to a Gaussian random variable. The consideration of odd-occupancy process was inspired by Spitzer, [48] to modeling the on/off events for lightbulbs. It is known that randomized odd-occupancy processes scale to a fractional Brownian motion with Hurst index $H\in(0,1/2)$ [10], and this type of limit theorems has been extended to stable processes indexed general metric spaces [9, 14]. The odd-occupancy mechanism is reflected in the stochastic-integral representations (see (2.16) below), and also in our representation for log-correlated bi-fractional Brownian motions.
The Karlin model with $H=0$ has been recently investigated by Iksanov et al., [21] and Iksanov and Kotelnikova, [22]. Limit theorems have been established in this regime with new Gaussian processes in the limit in the aforementioned papers. Their representation is not immediately recognizable to be related to ours and we provide some details in Appendix A. It should be expected that from their model a counterpart of our Theorem 3.1 can be established (they examined $j$-occupancy processes instead of odd-occupancy processes).
@@
$\log{\frac{\sqrt{t}+\sqrt{s}}{\sqrt{t}-\sqrt{s}}},\quad 0\leq s<t<\infty.$
Their limit theorem is in the same spirit as Neuman and Rosenbaum, [37].
Lifshits and Volkova, [32] also considered a limit theorem for $G^{H,K}$ as $K\downarrow 0$, although a Lamperti’s transformation is first applied to $G^{H,K}$ so that it is transformed to a stationary process before the limit is taken. A generalized process again arises in the limit.
Another example of a fractional Brownian motion with Hurst index $H=0$ has been seen in the studies of random matrix theory [15, 29]. The corresponding example is a centered self-similar Gaussian process with stationary increments.
@@ Acknowledgements
Y.W. is grateful to Mikhail Lifshits for several stimulating discussions on an earlier version of the paper, which lead to an enhancement of the results and a major revision of the paper. Y.W. thanks also Alexander Iksanov, Joseph Najnudel, and Yi Shen for several discussions. Y.W. was partially supported by Simons Foundation (MP-TSM-00002359), and a Taft Center Fellowship (2024–2025) from Taft Research Center at University of Cincinnati.
@@ 2. A stochastic-integral representation
In this and the next sections we consider ${\mathsf{M}}=[0,\infty)$, ${\mathsf{d}}(x,y)=|x-y|$. In most of this section we focus $H=1/2$ in the parameter of the bi-fractional Brownian motion except in Section 2.2.
Our reference for stochastic integrals with respect to stable random measures is Samorodnitsky and Taqqu, [45]: if $(E,{\mathcal{E}},\mu)$ is a measure space, $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on $(E,{\mathcal{E}})$ with control measure $\mu$, then the stochastic integral $\int f{\rm d}M_{\alpha}$ with control measure $\mu$ is defined as a symmetric $\alpha$-stable random variable with
@@
Let $(\Omega,{\mathcal{F}},\mathbb{P})$ be the by-default probability space that the random measures are defined on. Set $E=\mathbb{R}_{+}\times\Omega^{\prime}$ where $(\Omega^{\prime},{\mathcal{F}}^{\prime},\mathbb{P}^{\prime})$ is another probability space on which a standard Poisson process $N^{\prime}$ is defined, and $M_{2}$ is a Gaussian random measure on $\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$. Write
@@
We introduce
@@
and
(2.3) ${\mathcal{F}}_{0}:=\left\{f\in L^{\infty}([0,\infty)):\int_{1}^{\infty}|f(s)|\log s{\rm d}s<\infty\right\}.$
@@
The collection $\{G(f)\}_{f\in{\mathcal{F}}_{0}}$ is a family of centered Gaussian random variables with covariance function
${\rm{Cov}}(G(f),G(g))=\int_{0}^{\infty}\int_{0}^{\infty}f(s)g(t)\log\frac{s+t}{|s-t|}{\rm d}s{\rm d}t,\quad\mbox{ for all }f,g\in{\mathcal{F}}_{0},$
that also satisfies
(2.4) $G(a_{1}f_{1}+a_{2}f_{2})=a_{1}G(f_{1})+a_{2}G(f_{2})\mbox{ almost surely, for all }a_{1},a_{2}\in\mathbb{R},f_{1},f_{2}\in{\mathcal{F}}_{0}.$
@@
We claim that
@@
The first equality follows from the property of stochastic integrals. Next we show the expression on the right-hand side of (2.5) is finite. We prove $\int_{0}^{\infty}\int_{s}^{\infty}|f(s)f(t)\log((s+t)/(t-s))|{\rm d}t{\rm d}s<\infty$. To see this, we observe
$\int_{0}^{\infty}\int_{s+1}^{\infty}|f(s)f(t)|\left|\log{\frac{s+t}{t-s}}\right|{\rm d}t{\rm d}s\leq\int_{0}^{\infty}|f(s)|\int_{s+1}^{\infty}|f(t)|\log(4t){\rm d}t{\rm d}s<\infty,$
$\int_{0}^{\infty}\int_{s}^{s+1}\left|f(s)f(t)\log\frac{1}{t-s}\right|{\rm d}t{\rm d}s\leq{\left\|f\right\|}_{\infty}\int_{0}^{\infty}|f(s)|{\rm d}s\int_{0}^{1}|\log t|{\rm d}t<\infty,$
$\int_{1}^{\infty}\int_{s}^{s+1}\left|f(s)f(t)\right|\log(s+t){\rm d}t{\rm d}s\leq\int_{1}^{\infty}\int_{s}^{s+1}|f(s)f(t)|\log(2t){\rm d}t{\rm d}s<\infty,$
and
$\int_{0}^{1}\int_{s}^{s+1}\left|f(s)f(t)\log(s+t)\right|{\rm d}t{\rm d}s$ $\leq{\left\|f\right\|}_{\infty}^{2}\int_{0}^{1}\int_{s}^{s+1}\left|\log(s+t)\right|{\rm d}t{\rm d}s$
$\leq{\left\|f\right\|}_{\infty}^{2}\int_{0}^{1}|\log s|+\log(1+s){\rm d}s<\infty.$
It remains to prove the equality in (2.5), which follows from a careful application of Fubini’s theorem. Indeed, for all $0\leq s<t$,
@@
Here in the first step we applied Fubini’s theorem to interchange the two inner integrals (in this step when verifying the condition in Fubini’s theorem for each $r>0$, we bounded $|{{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{r}(s)||{{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)|\leq 1$ and used the assumption ${\left\|f\right\|}_{1}<\infty$), and in the second step we applied (2.6). In the last expression above, we can interchange the two integrals again by Fubini’s theorem, which, in combination of Frullani integral
(2.7) $\int_{0}^{\infty}\frac{e^{-r(t-s)}-e^{-r(t+s)}}{r}{\rm d}r=\log{\frac{t+s}{t-s}},\quad 0\leq s<t,$
leads to the equality in (2.5) (in this step we also need the fact that the right-hand side of (2.5) is finite, which we have shown earlier).
@@
where we follow the same notations for stochastic integrals earlier. Note that by stochastic Fubini theorem [38, Proposition 5.13.1], we have the representation
(2.9) $G^{(\epsilon)}(f)=\int_{0}^{\infty}\widetilde{G}^{(\epsilon)}_{t}f(t){\rm d}t,\mbox{ almost surely,}$
@@ 2.1. An extension to a generalized stable process
$G_{\alpha}(f):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}tM_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),$
where $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on $0,\infty)\times\Omega^{\prime}$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$. So $G(f)=G_{2}(f)$ and we again have $G_{\alpha}(a_{1}f_{1}+a_{2}f_{2})=a_{1}G_{\alpha}(f_{1})+a_{2}G_{\alpha}(f_{2})$ almost surely for $a_{1},a_{2}\in\mathbb{R},f_{1},f_{2}$ from a suitably chosen class of functions. Recall that $M_{\alpha}$ is a symmetric $\alpha$-stable random measure and the corresponding stochastic integrals are random variables characterized by (2.1). Formally, we have
(2.11) ${\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right)=\exp\left(-2\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}\right).$
Note that when $\alpha\in(0,2)$ it is not immediately clear what is the exact class of functions so that the right-hand side above is finite. Recall ${\mathcal{F}}_{0}$ in (2.3) and introduce
(2.12) ${\mathcal{F}}_{\delta}:=\left\{f\in L^{\infty}([0,\infty)):\int_{1}^{\infty}s^{\delta}|f(s)|{\rm d}s<\infty\right\},\delta>0.$
In particular, $f\in{\mathcal{F}}_{\delta}$ for any $\delta\geq 0$ implies that ${\left\|f\right\|}_{1}<\infty$.
@@
Assume one of the following:
  1. (i)
    $\alpha\in(0,2],f\in{\mathcal{F}}_{\delta}$ for some $\delta>0$, or
  2. (ii)
    $\alpha\in(1,2],f\in{\mathcal{F}}_{0}$.
Then, $G_{\alpha}(f)$ is a symmetric $\alpha$-stable random variable with characteristic function (2.11) (which is finite).
@@
(2.13) $\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left|\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}{\rm d}\mathbb{P}^{\prime}\\ =\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}<\infty.$
Without loss of generality, assuming $f\geq 0$. Introduce
$\Psi_{f}(r):=\int_{0}^{\infty}\int_{s}^{\infty}f(s)f(t)\left(e^{-r(t-s)}-e^{-r(t+s)}\right){\rm d}t{\rm d}s.$
We shall show
(2.14) $\Psi_{f}(r)\leq\begin{cases}C(r^{\delta}\wedge r^{-1}),&\mbox{ if }f\in{\mathcal{F}}_{\delta},\delta>0,\\ C(|\log r|^{-2}\wedge r^{-1}),&\mbox{ if }f\in{\mathcal{F}}_{0},\end{cases}\quad\mbox{ for all }r>0,$
where the constant $C$ depends on $f$ but not $r$. Then, by Hölder’s inequality, we have
${\mathbb{E}}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}$ $\leq\left({\mathbb{E}}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{2}\right)^{\alpha/2},$
$=\left(2\Psi_{f}(r)\right)^{\alpha/2},$
and hence (2.13) becomes $\int_{0}^{\infty}\Psi_{f}(r)^{\alpha/2}r^{-1}{\rm d}r<\infty$, which now follows from (2.14).
It remains to prove (2.14). First, for all $r>0$ we have
Next, assuming $f\in{\mathcal{F}}_{\delta}$ with $\delta>0$ first. By a change of variables write first
and
We have proved (2.14) with $f\in{\mathcal{F}}_{\delta},\delta>0$.
Now assume $f\in{\mathcal{F}}_{0}$. This time, fix $\gamma\in(0,1)$. We have,
$\int_{0}^{(1/r)^{\gamma}}\int_{0}^{\infty}f(s)f(s+t)e^{-rt}(1-e^{-2rs}){\rm d}t{\rm d}s$ $\leq 2r\int_{0}^{(1/r)^{\gamma}}sf(s)\int_{s}^{\infty}f(t){\rm d}t{\rm d}s$
$\leq 2r^{1-\gamma}{\left\|f\right\|}_{1}^{2},$
and, for $r<1/2$,
$\int_{(1/r)^{\gamma}}^{\infty}f(s)\int_{s}^{\infty}f(t){\rm d}t{\rm d}s$ $\leq\int_{(1/r)^{\gamma}}^{\infty}\frac{f(s)}{\log s}\int_{s}^{\infty}f(t)\log t{\rm d}t{\rm d}s$
$\leq\frac{1}{\gamma^{2}\log^{2}r}\int_{1/r}^{\infty}f(s)\log s{\rm d}s\int_{s}^{\infty}f(t)\log t{\rm d}t\leq C\frac{1}{\log^{2}r}.$
So for $r<1/2$ we have $\Psi_{f}(r)<C/\log^{2}r$. By possibly increasing $C$ we have proved (2.14) when $f\in{\mathcal{F}}_{0}$. ∎
2.2. A representation for the bi-fractional Brownian motion with $H\in(0,1/2)$
In the previous discussions in this section, we essentially worked with $G^{1/2,2H}$, of which the stochastic-integral representations have been known. Here, we provide a new stochastic-integral representation for $G^{H,K}$ with $H\in(0,1/2)$ (to be compared with the case $H=1/2$, which we recall in (2.17) below). Consider